A THIN HALF RING OF RADIUS 20cm IS UNIFORMLY CHARGE WITH A TOTAL CHARGE Q=0.7nC.THE MAGNITUDE OF ELECTRIC FIELD AT THE CENTRE OF THE HALF RING IS

Grade:12th pass

1 Answers

Vikas TU
7 years ago
We would need to calculate the electric field in both direction that is x and y respeectively.
Therefore for x direcn.,
dE = k’*k*dl*cos(thetha)/r^2
where k’ be the charge densite over the semi arc.
Now integerate both sides from theta 0 to pi.
O integerating we get,
Ex = (k*k’/r)(sinpi – sin0) = 0
Similarly in y direcn. we get,
dEy= k’*k*dl*cos(thetha)/r^2
Integerating both sides,
Ey = 2*K*K’/r
 
Net electric field => E = Ey => 2KK’/r
where K = 1/4pie
and K’ = 0.7 x 10^-9/pi*0.2.